# A Foundation Graph Model

### v2

@@ Abstract
The principal benefit of unsupervised graph representation learning is that a pre-trained model can be fine-tuned where data or labels are scarce. Existing approaches are domain specific, maintaining consistent node and edge attributes across the pre-training and target datasets. This precludes transfer to other domains. A model capable of positive transfer on arbitrary tasks and domains would represent the first foundation graph model.
In this work we use adversarial contrastive learning to present FoToM, a graph pre-training method based on node and edge feature exclusion. We use FoToM to pre-train models over multiple graph domains, producing the first foundation graph models. We demonstrate positive transfer on evaluation datasets from multiple domains, including domains not present in pre-training data. On all datasets performance is at worst on-par and on 76% significantly better than a supervised baseline, with an 8 to 40% reduction in error at 95% confidence. Contrary to other research, pre-training on a dataset with the target domain excluded leads us to better performance than pre-training on a dataset from only the target domain. The multi-domain model at worst, matches, and on 56% of tasks, significantly outperforms single-domain ($P\leq 0.01$). These results include when node labels are used in evaluation, where performance is consistently superior to single-domain or non-pre-trained models. Notably, FoToM benefits scenarios in both large or scarce data regimes for the target domains.
[…]
v#1
alexander.davies@bristol.ac.uk
${}^{1}$ School of Computer Science
${}^{2}$ School of Engineering Mathematics and Technology
University of Bristol, UK
@@ 1 Introduction
Graph learning has received considerable attention in recent years (Zhou et al., 2020). Despite the success of foundation language models, the question of how to train a foundation graph model remains open (Shirzad et al., 2022). Foundational language models have been trained through self-supervised representation learning. The success of these models is largely attributed to scaling up the size of data used for pre-training. Recent works for pre-trained graph models, where data-scaling is limited to just one domain, are yet to adopt this approach. The principle obstacle has been maintaining consistent node and edge features across the pre-training and downstream datasets. In order to develop graph foundation models, methods to scale graph data to multiple domains must be explored.
Graphs are a challenging form of data to process, given their high dimensionality. Methods normally applied to continuous or categorical data-types cannot be easily transferred to the graph domain. As such much work has been conducted on forming useful representations of graphs which are more easily used by well explored methods. Khoshraftar et al. (2022), Hamilton et al. (2017) and Chen et al. (2020) provide surveys on graph representation learning in general. These and other works focus on in-domain learning, where a single model is pre-trained and fine-tuned on data from a single domain. The principle motivation for this approach is maintaining consistent node and edge features between pre-training and downstream tasks. To the best of our knowledge there have been no works proposing to learn representations for multiple domains of graphs.
@@
In this work we present FoToM (Foundation Topology Models), a generalised pre-training method for graph models, which relies on excluding node and edge features. We create a multi-domain dataset, and using FoToM through contrastive learning, produce pre-trained models with consistently positive transfer on multiple domains. Under recent definitions these are the first foundation graph models (Bommasani et al., 2021; Liu et al., 2023; Zhang et al., 2023).
1.1 Hypotheses
We take the definition of a ‘foundation model’ from Bommasani et al. (2021), a definition since used by broad sources including world governments (Competition and Markets Authority, 2023), as “a type of AI technology …can be adapted to a wide range of tasks and operations”. Our interpretation given this definition is a single model that can be transferred with broadly positive performance increases on arbitrary graph tasks. From this, we present our two hypotheses:
Hypothesis 1 (H1):
There are useful generic structures within graphs, excluding node and edge features, that can be applied across domains.
Hypothesis 2 (H2):
@@ 1.1 Hypotheses
With the exclusion of node and edge features out-of-domain pre-training can have greater benefits than in-domain pre-training.
Figure 1: Current graph pre-training is domain-specific, with transfer only to near-domain samples, in order to maintain consistency for node and edge features.
@@
Contribution Our contribution, following these hypotheses, is a method for generalised pre-training of graph models. Our method, FoToM, relies on the exclusion of node and edge features during pre-training. In the interest of simplicity we use only node and edge labels as features. These are excluded during pre-training. We evaluate our method using graph constrastive learning with both random and adversarial augmentations. Through this evaluation we present the following findings:
@@
    Pre-training on varied domains, under the exclusion of node and edge features, leads to positive transfer.
@@
    Pre-training in this manner is more beneficial from varied non-target domains than from the target domain.
@@
Both of these findings go beyond the current consensus of graph pre-training, opening broad avenues of research and application. In particular the second finding directly opposes the current single-domain training paradigm.
Through the first finding we find that under fine-tuning results at worst on-par, and on 76% of tasks significantly better, than a non-pretrained baseline. At 95% confidence we observe a reduction in error of 8 to 40%. For the second finding results are similar. On molecular benchmarks we find that a model pre-trained on non-molecules, compared to pre-trained only on molecules, is at worst on-par. On 56% of tasks the non-molecule model performs significantly better, with 1 to 21% comparative decrease in error at 95% confidence.
2 Background
Here we define the notation and concepts necessary to understand the rest of this work. We then delineate the broad state of graph representation learning and the current obstacles to foundation graph models.
Representation Learning Let the input random variable $X\in\mathcal{X}$ represent data points in the input space $\mathcal{X}$. Similarly, the output random variable $Y\in\mathcal{Y}$ corresponds to the output space $\mathcal{Y}$, which is often the target space. Given their joint distribution $p(X,Y)$, the relevant information is defined as the mutual information $I(X;Y)$. Let $F$ denote a mapping $X\rightarrow\hat{X}$; finding an optimal representation $\hat{X}$ is formulated as minimising the following Lagrangian (Tishby and Zaslavsky, 2015):
$L[p(\hat{x}|x)]=I(X;\hat{X})-\beta I(\hat{X};Y)$ (1)
For unsupervised representation learning, the function $F$ is learnt without supervision on $Y$.
Contrastive Learning It is possible to make progress in minimising Equation 1 via contrastive learning (Suresh et al., 2021). Let $\mathcal{T}$ denote a set of augmentation functions where $t\in\mathcal{T}:X\rightarrow X^{\prime}$. Intuitively these functions make small alterations to the input sample. An encoder $F(\cdot)$ maps between $X$ and $\hat{X}$. From this encoding a projection head $g(\cdot)$ maps from $\hat{X}$ and $Z$. A sample $X$ undergoes two distinct augmentations, $X\rightarrow X^{\prime},X^{\prime\prime}$, and in combination the encoder and projection head are trained to maximize agreement between $Z$ for both augmented samples using a contrastive loss. After training is completed the encoder $F(\cdot)$ and representation $\hat{X}$ are used for for downstream tasks.
Graphs Let $G=(V,E)$, an un-attributed graph (topology) with node set $V=\{v_{1},v_{2},...,v_{n}\}$ and edges between them $E=\{(v_{1},v_{3}),(v_{2},v_{4}),...\}$. Let $n=|V|$ and $m=|E|$ denote the number of nodes and edges respectively. Let $X_{V}\in\mathbb{R}^{n\times d_{V}}$ denote the initial node feature matrix, where each node $i\in V$ has numeric vector node features of dimension $d_{V}$. Similarly edge features are denoted $X_{E}\in\mathbb{R}^{m\times d_{E}}$. With the inclusion of features and a graph-level label $y$ a graph is $\underline{G}=(V,E,X_{V},X_{E},y)$. As an example a molecule might have $V$ as atoms, $E$ as bonds between these atoms, $X_{V}$ atom labels and calculated features, $X_{E}$ bond types and counts, and $y$ a molecular property such as solubility.
2.1 Related Work
@@
Graph Representation Learning Hamilton (2020) provides a comprehensive summary of methods for Graph Representation Learning (GRL). Traditional methods to encode graphs as vectors include graph statistics such as centrality measures and clustering coefficients, but also kernel methods such as the Weisfeiler-Lehman (WL) kernel (Schweitzer PASCAL et al., 2011). Expressive Graph Neural Networks (GNNs) models, generally through parameterised message-passing, can produce powerful representations of graphs. A number of GNN architectures have been proposed with the Graph Isomorphism Network (GIN) being the most expressive (Zhou et al., 2020).
@@
Graph Contrastive Learning Multiple studies exist for graph contrastive learning (Chu et al., 2021; Hafidi et al., 2022; Hassani and Khasahmadi, 2020; Hu et al., 2020; Li et al., 2021; You et al., 2020; Zhu et al., 2021). The typical graph contrastive learning formulation, GraphCL, is put forwards by (You et al., 2020). They propose or describe four different graph augmentation strategies: 1) Node dropping, in which a subset of nodes are removed. 2) Edge perturbation, in which edges are added or removed. 3) Attribute masking, in which some subset of features are removed. 4) Subgraph sampling, in which subgraphs are used as augmentations. In GraphCL work and many subsequent works, such augmentations are performed randomly using fixed parameters, ie ‘drop 20% of nodes’.
@@
Adverarial Graph Contrastive Learning It has been shown that the quality of augmentations is critical for GCL (Zhu et al., 2021; You et al., 2020). To maximise expressivity in these augmentations adversarial strategies have been proposed (Hassani and Khasahmadi, 2020; Zhao et al., 2023). Suresh et al. (2021) propose Adversarial-Graph Contrastive Learning (AD-GCL), with their method training both an encoder and a ‘view-learner’. The idea is to perform the standard GCL protocol with weight updates to a parameterised set of augmenting functions $\mathcal{T}$ through the contrastive loss. Here $\mathcal{T}$ learns augmentations that minimise mutual information between the source and augmented graphs through edge dropping. The encoder, adversarially to this, learns to maximise shared information between the respective augmentations. The updates for $\mathcal{T}$ are regularised such that augmentations are not entirely destructive.
Graph Foundation Models Whilst improvements have been made towards contrastive learning methods, little work has been completed in methodologies to combine different domains of graph data effectively. Ubiquitous in other fields, primarily text and images, is pre-training on a large amalgamated dataset of samples. Archetypal to this approach is ImageNet (Russakovsky et al., 2015), a massive combined dataset of images that aims to span the full space of realistic or likely images. Previous authors have argued, but not shown, that an ImageNet equivalent for graphs is not feasible (Shirzad et al., 2022). Such an approach has been discussed in recent literature by Zhang et al. (2023). In this work the authors argue that the ‘extreme diversity’ present in graph data make producing a graph foundation model for multiple graph domains ‘exceedingly challenging, if not unlikely’. The assumption in this and prior works has been that the presence of node and edge attributes, with greatly varied meaning between domains, precludes generalised pre-training on an amalgamated dataset. We argue that this commitment to consistent features between pre-training and downstream tasks is the principle obstacle to the scale required for foundation models.
3 Method
Here we describe our method for evaluating FoToM, a method for generalised pre-training of graph models. Using FoToM node and edge features are excluded during pre-training This means an encoder is pre-trained on only the topology $G$, without the features or labels present in $\underline{G}$. This treatment of graphs $G$ as abstract structures, instead of specific examples, allows creation of a multi-domain amalgamated dataset. Recent works view the creation of such a graph dataset, given node and edge features, as unfeasible (Shirzad et al., 2022; Liu et al., 2023; Zhang et al., 2023).
3.1 Excluding Features
Figure 2: The Chem model is pre-trained only on molecular data, meaning in-domain transfer onto molecular datasets. The Social model is pre-trained only on non-molecular data, meaning out-of-domain transfer onto molecular datasets.
@@ 3.1 Excluding Features
Establishing Hypothesis 1 requires producing a model that has learnt useful representations of graphs in general. Our first step to achieving this is to construct a dataset that covers a reasonable breadth of real-world graphs. Constructing a dataset of real-world graphs is itself a reasonable narrowing of scope, but as the possible space of graphs at each number of nodes is so massive, constraining ourselves to a large and highly varied set of graphs is a reasonable step. Such a dataset should contain graphs that present a large diversity of multi-node patterns. Our approach is simple: include graphs from many domains, and assume that there is adequate coverage of the multi-node patterns that occur in real-world graphs. The same approach has essentially been taken in the domains of text (Brown et al., 2020) and images (Rombach et al., 2022), where enormous numbers of data samples are scraped from the web for generative pre-training. This relies on the assumptions that a) this dataset adequately covers the relevant space, and b) a single model can learn such a broad space. This assumption, as in other fields, is evaluated through downstream tasks on seen and unseen domains.
@@
We formulate our test for Hypothesis 1 as beating a simple supervised baseline. Said baseline is a model with the same architecture as the pre-trained model but without pre-training. Both models are then trained, or fine-tuned for the pre-trained model, on the same downstream datasets under the same conditions. On each downstream task, should the pre-trained model match or out-perform the supervised model, this represents positive transfer. Positive transfer on a task indicates that at least some of the learning during pre-training was beneficial to subsequent learning during fine-tuning. A positive transfer rate across datasets of more than 50% indicates that in-general transfer is positive, thus meeting the minimum criteria for Hypothesis 1. A high positive transfer rate would strongly establish Hypothesis 1.
3.2 Data Subsets
In order to evaluate Hypothesis 2, that domain diversity brings greater benefits than samples from the same domain, we need to produce models trained under different subsets of our data. The first model (Domain Model) is trained on a dataset from a single domain, as is the current status-quo. Said single domain matches that of the downstream target dataset(s). The second (Diverse Model) is instead trained on the rest of the data while excluding the target domain. The total size of the pre-training datasets for each model should match for fair comparison. By pre-training both models under the same controlled conditions (hyperparameters etc.), then evaluating their performance under fine-tuning, we can directly compare pre-training on a single domain to pre-training on multiple domains. A simple diagram for this process, specific to the experiments detailed in the next section, is presented in Figure 2.
In order to evaluate Hypothesis 2 we can apply both models on the given target domain. Verifying that Hypothesis 2 holds requires that positive transfer rates should be higher from the Diverse Model than from the Domain Model. The minimum threshold here is on-par positive transfer rates from each model. General positive transfer from the Diverse Model, and general negative transfer from the Domain Model, would constitute a clear result for Hypothesis 2.

### v3

# Towards Generalised Pre-Training of Graph Models

@@ Abstract
The principal benefit of unsupervised representation learning is that a pre-trained model can be fine-tuned where data or labels are scarce. Existing approaches for graph representation learning are domain specific, maintaining consistent node and edge features across the pre-training and target datasets. This has precluded transfer to multiple domains.
In this work we present Topology Only Pre-Training, a graph pre-training method based on node and edge feature exclusion. Separating graph learning into two stages, topology and features, we use contrastive learning to pre-train models over multiple domains. These models show positive transfer on evaluation datasets from multiple domains, including domains not present in pre-training data. On 75% of experiments, ToP models perform significantly ($P\leq 0.01$) better than a supervised baseline. These results include when node and edge features are used in evaluation, where performance is significantly better on 85.7% of tasks compared to single-domain or non-pre-trained models. We further show that out-of-domain topologies can produce more useful pre-training than in-domain. We show better transfer from non-molecule pre-training, compared to molecule pre-training, on 79% of molecular benchmarks.
@@ 1 Introduction
Much work has been conducted on forming useful representations of graphs which are more easily used by well-explored methods. Khoshraftar et al. [15], Hamilton et al. [12] and Chen et al. [4] survey graph representation learning in general. Such representation learning, after transfer and fine-tuning, allows greater performance in areas of data or label scarcity.
Whilst there is a good deal of research into graph representation learning methods, there is little work on methodologies to combine different domains of graph data effectively. Ubiquitous in other fields, primarily text and images, is pre-training on a large amalgamated dataset of samples, aiming to span the full or realistic space of the data [28]. The main obstacle to such an approach for graphs is consistency in node and edge features, as graph data presents a large diversity in possible feature-sets.
Given this variety of features between domains current research argues, but does not show, that multi-domain pre-training on graphs is not feasible [21, 31, 39]. There are very few works attempting learning across multiple domains of graphs [19]. This excludes areas of data or label-sparsity from the benefits of pre-training. Multi-domain pre-training will allow learning on domains where data is too scarse to form pre-training datasets.
@@
Here we propose Topology Only Pre-Training (ToP), a generalised pre-training method for graph models, under which pre-training is conducted with node and edge features excluded. We create a multi-domain dataset, and using ToP through adversarial contrastive learning, produce pre-trained models with consistently positive transfer across multiple domains and tasks. Transfer is also positive when node and edge features are included in downstream datasets during fine-tuning.
In this work we propose and investigate ToP as a pre-training method for graph data that is not limited to a given domain or task. The result of such pre-training is a single model that can be transferred with broadly positive performance increases on arbitrary graph tasks and domains. Experimentally we test two hypotheses:
Hypothesis 1: Excluding Features
Pre-training on multiple domains with node and edge features excluded can lead to consistent positive transfer compared to fully-supervised, task-specific models.
Hypothesis 2: Sample Diversity
@@
With the exclusion of node and edge features out-of-domain pre-training can have greater benefits than in-domain pre-training across task-relevant metrics.
@@
Findings We evaluate ToP using graph constrastive learning with both random and adversarial augmentations. Through this evaluation we contribute the following empirical findings:
@@
    Effective pre-training without features. Pre-training on varied domains, under the exclusion of node and edge features, leads to positive transfer. We find that under fine-tuning, results are at worst on-par, and on 76% of tasks significantly better, than a non-pretrained baseline.
@@
    Increased performance from domain diversity. Pre-training without features from varied non-target domains, instead of the target domain, can bring greater downstream performance benefits. On molecular benchmarks, we find that a model pre-trained on non-molecules, compared to pre-trained only on molecules, is at worst on-par. On 56% of tasks the non-molecule model performs significantly better, and never significantly worse, than the molecule model.
@@
Given these findings, under recent definitions, these may be among the first foundation graph models [2, 19, 21, 39]. Both of these findings go beyond the current consensus of graph pre-training, opening broad avenues of research and application. In particular our second finding runs contrary to the single-domain training paradigm.
Organisation In Section 2, we describe the notation and concepts necessary to understand our method and provide an overview of related works. In Section 3, we describe our generalised pre-training method ToP. In Section 4, we describe how we evaluate the hypotheses detailed above, then detail the experiments we perform to evaluate ToP and these hypotheses. In Section 5 we discuss the implications of the results from these experiments, as well as list the assumptions we make and threats to validity. In Section 6, we conclude with avenues for future research.
2 Background and Related Works
Here we describe the concepts necessary to understand the rest of this work.
Representation Learning Let $X\in\mathcal{X}$ represent a random variable $X$ in the data space $\mathcal{X}$, and $Y\in\mathcal{Y}$ the same for the output space. Representation learning aims, given their joint distribution $P(X,Y)$, to learn a mapping $X\rightarrow\hat{X}$ that maximises their mutual information $I(X;Y)$. In most contexts this means producing numerical vectors for input data that usefully describe that data.
Contrastive Learning Let $\mathcal{T}$ be set of small transformations called ‘augmentations’, where $t\in\mathcal{T}:X\rightarrow X^{\prime}$. These augmentations are generally designed to manipulate the input sample in a way non-destructive to the ‘useful’ information it represents. Common choices for images, for example, are gray-scaling and rotations. Given an encoder’s mapping $F(X_{i})\rightarrow\hat{X_{i}}$, multiple augmentations of the same input sample should be as similar as possible.
@@ 2 Background and Related Works
Graph Representation Learning Hamilton [11] provides a comprehensive summary of methods for Graph Representation Learning (GRL). Traditional methods to encode graphs as vectors include graph statistics such as centrality measures and clustering coefficients, but also kernel methods such as the Weisfeiler-Lehman (WL) kernel [29]. Expressive Graph Neural Networks (GNNs) models, generally through parameterised message-passing, can produce powerful representations of graphs. A number of GNN architectures have been proposed with the Graph Isomorphism Network (GIN) being the most expressive [41] in representation learning contexts.
@@
Graph Contrastive Learning Multiple studies exist for graph contrastive learning (GCL) [5, 10, 13, 14, 18, 38, 42]. The typical graph contrastive learning formulation, Graph Contrastive Learning (GraphCL), is put forward by You et al. [38]. In GraphCL and many subsequent works, such augmentations are performed randomly using fixed parameters, i.e., ‘drop 20% of nodes’.
@@
Adversarial Graph Contrastive Learning It has been shown that the quality of augmentations is critical for graph contrastive learning [38, 42]. To maximise expressivity in these augmentations adversarial strategies have been proposed [13, 40]. Suresh et al. [32] propose Adversarial-Graph Contrastive Learning (AD-GCL), with their method training both an encoder and a ‘view-learner’ $\mathcal{T}_{\theta}$. AD-GCL performs the standard GCL protocol with weight updates to a parameterised ($\theta$) set of augmenting functions $\mathcal{T_{\theta}}$ through the contrastive loss. Here $\mathcal{T_{\theta}}$ learns augmentations that minimise mutual information between the source and augmented graphs through edge dropping. The encoder, adversarially to this, learns to maximise shared information between the respective augmentations. The updates for $\mathcal{T}$ are regularised such that augmentations are not entirely destructive.
One For All A work concurrent with ours is One-For-All (OFA) [19]. OFA employs a single model on graph and node classification tasks across domains, sidestepping the issues around node and edge features we’ve discussed by assuming that natural language descriptions of features are adequate, and using language model embeddings as features. OFA, then, requires that input graphs have textual features, or that existing features can be usefully described in natural language. Whereas OFA produces only classification models, ToP produces pre-trained models intended for downstream fine-tuning. As such ToP allows arbitrary downstream tasks to be conducted which are not possible with OFA, regression being an example.
3 ToP: Topology Only Pre-Training
Figure 1: The process for handling features with ToP. During pre-training, features are replaced with a single integer label for each node and edge (all identical), which are then passed to a single-layer MLP. This results in an identical input vector for each node. During transfer, a different input head is used, moving the original dimensions to the hidden dimensionality of the GNN encoder block. This allows arbitrary node and edge features to be included in transfer. Omitting the output head of the model allows transfer onto node and edge-level tasks.
One naive approach to generalised pre-training with varied feature sets included would be to pass features as-is, with padded vectors where feature sets are of a lower dimensionality. Here components would be shared across domain features, for example types of atoms and number of social followers in the same component, representing an extremely difficult space to learn. Alternatively one could reserve components of a long feature vector for each feature set, with padding tokens in components for other domains and feature sets. This would require a model to learn a massive span of features, and preclude movement to new feature sets, and again learning is unlikely to occur. Feature-management approaches, which move domain features to the same feature set [19], are in their infancy.
Given that pre-training while including features in their original form is not feasible, we view graph learning as a task occurring on both topologies $G$ and features on that topology $X$. From this view, pre-training without features is well-motivated, with models learning generalised information about graph structures before features are introduced. This information from $G$ is useful for downstream tasks, even while missing task-specific information, and in situations where data is abundant pre-training on both $X$ and $G$ will likely still lead to higher performance.
This is an abstraction away from feature-specific correlations, with abstraction a core quality of effective representation learners [16]. Pre-training without topologies, viewed as an initialisation, would provide a smoother loss landscape for downstream tasks [20]. In this case a wider range of topologies could represent more useful information than a narrow range closer to a downstream domain.
Topology Only Pre-Training (ToP) is a method for generalised pre-training of graph models. Under ToP pre-training is conducted only on topologies, with node and edge features excluded or replaced with identical values. This use of only topologies $G$, instead of graphs with specific features $\underline{G}$, allows pre-training on a multi-domain dataset. Prior to this work, the use of such a graph dataset with mixed node and edge features was not considered feasible [21, 31, 39].
3.1 Hypothesis 1: Excluding Features
@@
Supporting Hypothesis 1 requires producing a model that has learnt useful representations of graphs in general. Our first step to achieving this is to construct a dataset that covers a variety of graph domains. Such a dataset should contain graphs that present a large diversity of multi-node patterns. Our approach is simple: include graphs from several domains and assume that there is adequate coverage of the multi-node patterns that occur in real-world graphs for a given set of downstream tasks. The same approach has been taken in the domains of text [3] and images [24], where enormous numbers of data samples are scraped from the web for generative pre-training. This relies on two assumptions 1) this dataset adequately covers the relevant space for downstream transfer, and 2) a single model can usefully represent such a space. These assumptions, as in other fields, are justified through performance metrics on downstream tasks on seen and unseen domains.
@@
We formulate our test for Hypothesis 1, that pre-training on graphs without node or edge features is beneficial, as beating a simple supervised baseline. Said baseline is a model with the same architecture as the pre-trained model but without pre-training. Both models are then trained, or fine-tuned for the pre-trained model, on the same downstream datasets under the same conditions. On each downstream task, should the pre-trained model match or out-perform the supervised model, this represents positive transfer. Positive transfer on a task indicates that at least some of the learning during pre-training was beneficial to subsequent learning during fine-tuning. A positive transfer rate across datasets of more than 50% indicates that in-general transfer is positive, thus meeting the minimum criteria for Hypothesis 1. A high positive transfer rate strongly supports Hypothesis 1.
3.2 Hypothesis 2: Sample Diversity
From Hypothesis 2 we expect that for a given task models trained on a more diverse pre-training domain can bring better transfer than a model trained on the target domain. We do not argue that this is always true, only that in some settings out-of-domain samples can be more useful in pre-training than in-domain. This has clear parallels to the performance-by-scale approach taken in other fields [28, 6, 23]. To this end we compare two models, the first trained on a dataset from a single domain, as is the current practice. This single domain matches that of the downstream target dataset(s). The second model is instead trained on the rest of the data while excluding the target domain.
The total size of the pre-training datasets for each model should match for fair comparison. By pre-training both models under the same controlled conditions (hyperparameters etc.), then evaluating their performance under fine-tuning, we directly compare pre-training on a single domain to pre-training on multiple domains. General positive transfer from the multi-domain model, in comparison to transfer from the in-domain model, would constitute a clear supporting evidence for Hypothesis 2.